The sum of two numbers is 10. Three times the smaller number is equal to twice the larger number. What are the numbes? (_,_) Please help.
Answer by scott8148(6628)

"The sum of two numbers is 10" ___ x + y = 10 ___ x = 10 - y

"Three times the smaller number is equal to twice the larger number" ___ 3x = 2y

substituting ___ 3(10 - y) = 2y

solve for y, then substitute back to find x